40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 794     Entry time: Mon Jun 16 17:52:37 2014
Author: Gabriele 
Type: DailyProgress 
Category:  
Subject: Some progress 

 I opened the chamber and spent half afternoon tweaking the fiber feed-through. This thing is extremely unreliable. 

At the end I was able to get back a reasonable level of power, even not the maximum.

I also checked that the blades are not rubbing anywhere and that the shadow sensors are close to mid range. Nevertheless, after closing again, I'm still unable to drive at 800 cts.

I started a set of measurements at 600 counts.

 

ELOG V3.1.3-